# Self-supervised Pre-training for Mapping of Archaeological Stone Wall in Historic Landscapes Using High-Resolution DEM Derivatives

### v1

@@ 6 Conclusions
In this study, we introduced DINO-CV, a cross-view self-supervised learning framework for mapping dry-stone walls in archaeologically significant and visually complex landscapes. Leveraging multi-view DEM derivatives of archaeological landscape, MHS and VAT, DINO-CV enables models to learn view-invariant and semantically rich features in self-supervised pre-training. Our method significantly improves stone wall mapping performance over both supervised and existing self-supervised baselines across ConvNet and ViT backbones. Under limited annotation, DINO-CV demonstrates strong label efficiency and generalization, making it highly suitable for large-scale cultural heritage mapping tasks.
@@
An inevitable limitation lies in the accuracy of wall annotations in the BudjBimArea dataset. Given the challenges in visually interpreting dense vegetation, shadowing, and subtle terrain features, some wall labels may suffer from spatial uncertainty or omission. Additionally, while the dataset focuses on a geographically and culturally rich region, its domain specificity may limit the direct transferability of results to other archaeological landscapes with different geomorphological characteristics or wall typologies.
@@
Future work could explore improving annotation quality of BudjBimArea dataset through semi-automated labeling strategies and integrating expert feedback loops. Additionally, extending the DINO-CV framework to support multi-modal data fusion, including hyperspectral imagery and spatial-temporal data, could potentially further enhance model’s mapping capability and robustness. Finally, applying DINO-CV to other cultural heritage sites worldwide for stone wall mapping could help assess its generalizability and scalability.
@@
Zexian Huang: Conceptualization, Writing – original draft, Writing – review & editing, Software, Methodology, Investigation. Mashnoon Islam: Conceptualization, Data curation. Brian Armstrong: Writing – review & editing. Kourosh Khoshelham: Writing – review & editing, Supervision. Martin Tomko: Writing – review & editing, Supervision.
@@
The source code and data support the findings of this study are available in following link: https://github.com/MLinArcheaomatics/BudjBimStoneWall-DINO-CV.
@@ Acknowledgements
We thank Gunditj Mirring Traditional Owners Corporation (GMTOC) for providing the research data and overseeing its usage in this study, their invaluable cooperation, support, and meticulous efforts have significantly contributed to the quality of our research.

### v3

@@ 6 Conclusions
In this study, we introduced DINO-CV, a cross-view self-supervised learning framework for mapping dry-stone walls in archaeologically significant and visually complex landscapes. Leveraging multiple DEM-derived views of archaeological landscape, MHS and VAT, DINO-CV enables models to learn view-invariant and semantically rich features in self-supervised pre-training. Our method significantly improves stone wall mapping performance over both supervised and existing self-supervised baselines across ConvNet and ViT backbones. Under limited annotation, DINO-CV demonstrates strong label efficiency and intra-site generalization, making it highly suitable for large-scale cultural heritage mapping tasks.
@@
An inevitable limitation lies in the accuracy of wall annotations in the BudjBimArea dataset. Given the challenges in visually interpreting dense vegetation, shadowing, and subtle terrain features, some wall labels may suffer from spatial uncertainty or omission. Additionally, while the dataset focuses on a geographically and culturally rich region, its domain specificity may limit the direct transferability of results to other archaeological landscapes with different geomorphic characteristics or wall typologies.
@@
Future work could further explore improving the annotation quality of the BudjBimArea dataset through semi-automated labeling strategies and integrating expert feedback loops. Additionally, extending self-supervised pre-training to cross-platform transfer learning on LiDAR data (UAV-airborne LiDAR) represents a promising research direction. For instance, while UAV-LiDAR enables fine-grained capture of low-relief archaeological features at local scales, acquiring reliable annotations remains challenging and resource-intensive. The cross-view pre-training strategy of DINO-CV, designed to extract invariant geometric features from multiple DEM derivatives, is theoretically well suited to bridge the domain gap between regional airborne LiDAR surveys and high-resolution UAV-LiDAR acquisitions. Transferring models pre-trained on extensive airborne datasets to UAV-based surveys could therefore mitigate annotation scarcity and enable high-resolution mapping without requiring extensive new labels. Finally, applying DINO-CV to additional cultural heritage landscapes worldwide would help further assess its generalizability and scalability across diverse geomorphic settings and stone wall typologies.
@@
Zexian Huang: Conceptualization, Writing – original draft, Writing – review & editing, Software, Methodology, Investigation. Mashnoon Islam: Conceptualization, Data curation. Brian Armstrong: Writing – review & editing. Billy Bell: Project administration, Supervision. Kourosh Khoshelham: Writing – review & editing, Supervision. Martin Tomko: Writing – review & editing, Supervision.
@@
Funding sources
This research was supported by the Australian Research Council (ARC) under grant SR200200227. The authors gratefully acknowledge the Department of Environment, Land, Water and Planning (DELWP), Victoria, for sourcing and providing access to the airborne LiDAR dataset. All data and derived results are owned by the Gunditj Mirring Traditional Owners Corporation (GMTOC), with research conducted in collaboration under an established memorandum of understanding (MoU).
Declaration of generative AI use
During the preparation of this work, the authors used ChatGPT (OpenAI, San Francisco, USA) to assist in improving language clarity, grammar, and readability of the manuscript. All content was carefully reviewed, verified, and edited by the authors to ensure accuracy, coherence, and that the final text reflects their original research, interpretation, and conclusions. The authors take full responsibility for the content of the published article.
@@
The source code and data supporting the findings of this study are available at: https://github.com/MLinArcheaomatics/BudjBimStoneWall-DINO-CV.
@@ Acknowledgements
We thank Gunditj Mirring Traditional Owners Corporation (GMTOC) for providing the research data and overseeing its use in this study. Their invaluable cooperation, support, and meticulous efforts significantly contributed to the quality of this research.